OpenStudy (anonymous):

definite integral from 1 to 2 of sqrt((x^2) - 1)/x. I recognize that if x = sec(t), then the function becomes tan(t)/sec(t), but where do I go from there? Thanks.

OpenStudy (dumbcow):

you have to replace dx with dt x = sec(t) dx = sec(t) tan(t) dt \[\int\limits \frac{\tan t}{\sec t} \sec t \tan t dt = \int\limits \tan^{2} t dt = \int\ sec^{2}t -1 = \tan t - t\]

OpenStudy (anonymous):

Thank you so much! How do I translate the 1 to 2 into the substituted integral?

OpenStudy (dumbcow):

well \[t = \sec^{-1}(x)\] \[\sec^{-1}(1) = 0\] \[\sec^{-1}(2) = \pi/3\]

OpenStudy (anonymous):

Ah, yes. That's right. I appreciated the help.

OpenStudy (dumbcow):

yw